OpenStudy (vishweshshrimali5):

If the curve \(y = x^4 + 4rx + 3s\) , where r and s are real numbers; lies entirely above the x - axis, then: (a) \(r^2 < s^3\) (b) \(r^2 > s^3\) (c) \(r^4 s^3\)

OpenStudy (vishweshshrimali5):

@ganeshie8 @Miracrown @mukushla

OpenStudy (vishweshshrimali5):

What I did was this: For the curve to lie entirely above the x-axis; its minimum value must be positive. So, using maxima - minima principle, \(y' = 4x^3 + 4r\) Putting y' = 0 \(4x^3 + 4r = 0\) \(\large{\implies x^3 = -r}\) \(\large{\implies x = -r^{\cfrac{1}{3}}}\) Now, \(\large{y'' = 12x^2}\) which will always be positive irrespective of value of x. Thus, there is minima at \(\large{x = -r^{\cfrac{1}{3}}}\). So, \(\large{y_{min} = r^{\cfrac{4}{3}} - 4r^{\cfrac{4}{3}} + 3s}\)

OpenStudy (vishweshshrimali5):

Now, \(y_{min} > 0\) Thus, \(\large{y_{min} = r^{\cfrac{4}{3}} - 4r^{\cfrac{4}{3}} + 3s > 0}\) \(\large{\implies-3r^{\cfrac{4}{3}} + 3s > 0}\) \(\large{\implies s > r^{\cfrac{4}{3}}}\) \(\large{\implies s^3 > r^4}\) Thus, option (c) should be correct. Any alternative method from your side ?

OpenStudy (vishweshshrimali5):

@mathslover

OpenStudy (vishweshshrimali5):

@dan815

OpenStudy (dan815):

to find s and r such that only complex roots exist

mathslover (mathslover):

That was such a great method.. @vishweshshrimali5 great application of maxima and minima. Awesome!

OpenStudy (dan815):

yeah ^ :)